# Denoising Diffusion Probabilistic Models for Coastal Inundation Forecasting

### v1

@@ 6.3 Baseline Comparison
These experiments highlight that existing spatiotemporal forecasting methods (i.e., DiffSTG and DCRNN) have moderate performance but poor computation scalability, making them infeasible to deploy in large spatiotemporal settings. On the other hand, temporal forecasting methods do not have good performance scalability, making them inadequate for large-scale spatiotemporal forecasting tasks. Diff-Flood has both computation and performance scalability, making it suitable for large-scale spatiotemporal forecasting tasks.
Finally, the number of parameters of the baseline models grows almost exponentially with increasing patch sizes ($D$), as observed from Figure 3. To investigate the reason, we analyzed the implementation of TimeGrad as an example and found that it uses a fully connected layer proportional to the number of variables. On the other hand, our model utilizes a UNet architecture that creates a low-dimensional latent representation of the patches. It then performs expensive operations (e.g., feed-forward, cross-attention) on the lower-dimensional latent space, allowing it to scale quite well with the number of variables.
6.4 Ablation Study
Figure 4: Bar-plots showing performance comparison of the four configurations of Diff-Flood on our test dataset of configuration $64\times 64\times 10$, in terms of two metrics. The height of the bars represent mean of the metric and the vertical black lines represent standard deviation. Lower metric value is better. Our proposed architecture of Diff-Flood is able to utilize elevation (ELEV) and temporal covariates (COV) individually, resulting in improved performance in terms of both of our two metrics. Best performance is achieved when both are used.
In this section, we perform an ablation study of the Diff-Flood method to investigate the effects of utilizing elevation matrix and temporal covariates, in addition to the inundation history. Specifically, given inundation history (INUN: ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$), elevation (ELEV: ${}_{k}\mathbf{s}$), and temporal covariates (COV: $\mathbf{z}_{t_{0}:t-1}$), we define four configurations based on how the context embedding is computed:
  1. 1.
    INUN: Only the historical inundation (${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$) used.
  2. 2.
    INUN, ELEV: We use historical inundation data ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$ and the elevation matrix ${}_{k}\mathbf{s}$.
  3. 3.
    INUN, COV: We use historical inundation data ${}_{k}\mathbf{x}^{0}_{t_{0}:t-1}$ and the temporal covariates $\mathbf{z}_{t_{0}:t-1}$.
  4. 4.
    INUN, ELEV, COV: We include all three components.
We ran five experiment runs for each configuration, where we trained and tested Diff-Flood on the Tidewatch dataset with $K=10$ and $D=64$. Figure 4 shows the performance of the four configurations in-terms of NACRPS and NRMSE. We observe that, configuration (INUN, COV) and (INUN, ELEV) are both better than the configuration (INUN) in terms of the two metrics. This implies that Diff-Flood is able to utilize both elevation and temporal covariates as context effectively. We also observe that, including the temporal covariates results in a greater improvement ($43\%,41\%$ improvement over INUN in terms of NACRPS and NRMSE, respectively) than including spatial context ($8\%,4\%$). However, best performance is achieved when both are utilized ($48\%,47\%$ improvement over INUN).
7 Discussion and Conclusion
Once Diff-Flood is trained, we can use it to answer inundation-scenario related queries relevant to policymakers. Example queries are as follows:
  • Query 1: What is the probability that the flooding level in an area $A$ will be above $d$ units within next $T$ hours? (Figure 5, left)
  • Query 2: Given a route, what is the probability that the route will not be flooded in the next $T$ hours? (Figure 5, right)
Such queries are of special interest to city officials; answer to these queries will help to better prepare and respond to emergencies. The key idea for answering such queries is that, we have a sampling method in the form a trained Diff-Flood model and we can use it to sample inundation scenarios from the distribution of possible scenarios. We use the sampled scenarios to calculate our desired probability value. In the supplementary materials, we provide a detailed algorithm for answering such queries using a trained Diff-Flood model.
Although Diff-Flood is developed to solve the task of coastal inundation forecasting, it is generalizable enough to be extended to other spatiotemporal forecasting tasks (e.g., occupancy detection, traffic forecasting). Diff-Flood supports conditioning on time-invariant spatial features, as well as, on temporal co-variates. In our future work, we plan to evaluate its performance in diverse spatiotemporal forecasting tasks.
To conclude, we have presented a probabilistic spatiotemporal forecasting method, Diff-Flood that takes both spatial and temporal context into account during its training and forecasting phase. We have performed a case study on the coastal inundation forecasting task for the Eastern Shore of Virginia. Through our experiments, we show that Diff-Flood outperforms existing forecasting methods in predicting coastal inundation. Additionally, the architecture of Diff-Flood allows it to scale to large forecasting tasks at fine spatial resolution, compared to current methods.
Appendix A Extended Related Work
@@
Hydrology Models Traditionally, inundation has been simulated through hydro-dynamical physics-based 1D/2D dual drainage models. By simulating water flow in water bodies and coupling them with 1D or 2D hydrodynamic models, these models can predict inundation in surfaces. For example, the 2D model SOBEK (Deltares, 2014) consists of a 1D network with information about the initial water volume of rivers and 2D rectangular computational cells. By coupling these two parts, the model uses the Saint-Venant flow equation and 2D shallow water equation to calculate discharge between cells and determine the water level at each cell of the Digital Elevation Model (DEM). TELEMAC-2D (Vu et al., 2015) also solves 2D shallow water equation to simulate water flow. However, its primary purpose is to understand flow through water bodies. Therefore, this makes this model more suitable for coastal regions, contrary to SOBEK which is more suitable for urban areas. Polymorphic models like SCHISM (Zhang et al., 2016) can capture water flow at very high resolution, down to even a few meters due to their hybrid grinding system and localized coordinate system. Despite these models’ ability to model inundation accurately at fine resolution, the computation time associated with them makes them infeasible for real-time forecasting.
@@
Regression-based Models Early approaches in finding surrogates for hydrology models involved applying various regression-based methods. (Chang et al., 2010) first applied K-means clustering to pre-process data to find inundation control points. The authors then applied linear regression or neural networks to predict inundation for grids depending on whether they are linearly correlated with the control points or not. (Bermúdez et al., 2019) explored non-parametric regression approach based on least square support vector machine (SVM) as surrogate for 2D hydrodynamic models. (Lee and Kim, 2021) created an offline rainfall-runoff-inundation database by simulating hydrology models over multiple scenarios. Then, they trained a logistic regression-based approach based on the database to estimate the parameters to predict the risk of flooding for each region. Their goal was then to use the trained regressor to generate near real-time prediction of inundation probability given the runoff to the grid from simulated and realistic rainfall scenario.
@@
Deep Learning (DL) Models Many DL models have been explored as possible surrogates for Physics Based Models (PBM). One research work (Pan et al., 2011) applied a combined approach consisting of principal component analysis (PCA) and a feed-forward neural network. Their approach involved inundation prediction at various locations in the next timestep by looking at a fixed window of rainfall history of $N$ different rainfall gauges. First, they applied PCA on these various rainfall gauge histories to reduce the dimensionality before using them as input to the feed-forward neural network. Since such neural networks are expected to fail where data availability is limited, another work (Xie et al., 2021) explored the potential of a hybrid feed-forward neural network-based approach where they used two block-based neural networks along with the traditional point-based neural network and found superiority of the hybrid approach over other in the data-sparse region. Consequently, convolutional neural network (CNN) was explored (Kabir et al., 2020) as a potential surrogate. After their model was trained to predict the output simulated from a 2D hydrology model, they used it to simulate two real-event flood scenarios in a region of UK. Since these models are not able to capture long-term temporal dependencies, researchers have recently explored (Roy et al., 2023) seq2seq LSTM model as surrogates to hydrology models. Such model was also applied in forecasting for multi-step future horizons.
@@
Diffusion Model TimeGrad (Rasul et al., 2021) was the pioneering work that proposed the use of diffusion models for time-series forecasting tasks. They used temporal co-variates and embedded historical data as conditions for the diffusion model to denoise the prediction for the next time step. CSDI Tashiro et al. (2021) does the same but instead of encoding with a Recurrent Neural Network (RNN) as done in TimeGrad, they leverage a transformer-based architecture. SSSDS4 Alcaraz and Strodthoff (2022) uses a S4 model Gu et al. (2021) to capture the encoding which is particularly useful in capturing long-term dependencies. Also, instead of performing diffusion along both temporal and input channel dimensions, they perform diffusion only along the temporal dimension. Different from these, $D^{3}$ (Li et al., 2022) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability. The State-of-the-Art Diffusion-based model DiffSTG (Wen et al., 2023) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively. However, both these methods scale poorly with the growth in number of variables.
@@ Appendix B Denoising Diffusion Model
[…]
@@
Here, we get (21) by applying Jensen’s inequality. We indirectly minimize the negative log-likelihood of the data, i.e., $-\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (22) . We can show that it is equal to:
@@
Ho et al. (2020) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form . Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$. Then, we have:
@@
$=\frac{1}{2\tilde{\beta}_{n}}||\tilde{\mu}_{n}(\mathbf{x}^{n},\mathbf{x}^{0})-\mu_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}+C$ (29)
@@
$=\frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\ \beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (32)
@@
$\mathop{\mathrm{argmin}}_{\theta}\ \mathbb{E}_{n\sim U[2,N]}\ \mathbb{E}_{q(\mathbf{x}^{n}|\mathbf{x}^{0})}\frac{1}{2\tilde{\beta}_{n}}\frac{\bar{\alpha}_{n-1}\beta_{n}^{2}}{(1-\bar{\alpha}_{n})^{2}}||\mathbf{x}^{0}-\hat{\mathbf{x}}_{\theta}(\mathbf{x}^{n},n)||_{2}^{2}$ (33)
@@
$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{v}$ (34)
@@
Here, $\mathbf{v}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$ and $\mathbf{v}={\mathbf{0}}$ when $n=1$. $\mu_{\theta}(\mathbf{x}^{n},n)$ and $\Sigma_{\theta}$ are computed using (31) and (30), respectively.
Appendix C Inundation Scenario Queries

### v2

@@ 4    $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$
These experiment results highlight that existing spatiotemporal forecasting methods have either poor computational scalability or performance scalability, making it challenging to apply them to large-scale high-resolution spatiotemporal forecasting tasks.  Diff-Sparse offers both computation and performance scalability, making it suitable for large-scale high-resolution spatiotemporal forecasting tasks under sparse observations.
Figure 2: Bar-plots showing Diff-Sparse performance at varying sparsity levels in two patch configurations. Ten experiment runs were performed in each setting; mean and standard deviation of the performance metrics are shown.
Ablation Study
In this section, we perform ablation studies on Diff-Sparse to understand the importance of different context data. First, we investigate the predictive performance of Diff-Sparse at various sparsity levels. We trained and tested Diff-Sparse at four different sparsity levels ($0\%$, $50\%$$95\%$, and $100\%$) for the patch configurations $(64^{2},10)$ and $(64^{2},20)$. Sparsity level of $100\%$ means no sensor data is available. Figure 2 shows the performance of Diff-Sparse in these settings, in terms of our two performance metrics. We observe that in both patch configurations, Diff-Sparse has the best performance at $0\%$ sparsity level in terms of both metrics. As the sparsity level increases, performance of Diff-Sparse degrades; the worst performance observed at $100\%$ sparsity level. It implies that, having (more) sensors, i.e., observed inundation values in the context window, helps Diff-Sparse to make better predictions.
Figure 3: Ablation study at varying sparsity levels for two patch configurations. Ten experiment runs were performed in each setting; mean value of the performance metrics are shown using bar-plots.
Next, we investigate the utility of elevation data and temporal co-variates as additional context. Given sparse inundation history (INUN: $_{k}\mathbf{y}^{0}_{t_{0}:t-1}$), elevation (ELEV: $\mathbf{s}_{k}$), and temporal co-variates (COV: $\mathbf{z}_{t_{0}:t-1}$), we examine four configurations of Diff-Sparse based on how the context embedding is computed: (1) INUN: only sparse inundation values used, (2) INUN, ELEV: sparse inundation values and elevation data used, (3) INUN, COV: sparse inundation values and temporal co-variates used, (4) INUN, ELEV, COV: all three components used. Figure 3 shows the performance of these four settings at three different sparsity levels ($0\%$, $50\%$, and $95\%$) for the patch configurations $(64^{2},10)$ and $(64^{2},20)$. We see that in all sparsity levels, both (INUN, COV) and (INUN, ELEV) performs better than INUN; meaning both elevation data and temporal co-variates are individually helpful in making better forecasts. At low sparsity levels ($0\%$, $50\%$), the setting (INUN, COV) performs better than (INUN, ELEV), indicating the higher utility of temporal co-variates than elevation at low sparsity levels. However, at high sparsity level ($95\%$), we see the opposite; (INUN, ELEV) performs better than (INUN, COV). It means at high sparsity level, elevation data becomes more useful compared to temporal co-variates. In all sparsity levels, using all three components (i.e., INUN, COV, ELEV) yields the best result.
Conclusion
In this paper, we have addressed the problem of high resolution probabilistic coastal inundation forecasting from sparse sensor observations. To solve it, we first formulated it as a problem of learning a conditional distribution function. We then presented Diff-Sparse, a scalable spatiotemporal forecasting method that utilizes sparse sensor observations, elevation data, and temporal covariates to make probabilistic forecasts. Diff-Sparse is trained using data generated by a physics-based hydro-dynamical model. A novel masking strategy is employed to tackle the challenge of making forecasts based on sparse sensor observations during inference. Moreover, due to our robust training strategy, Diff-Sparse can make forecasts based on different sensor placement configurations without requiring retraining from scratch. Through our experiments, we demonstrated that Diff-Sparse outperforms existing forecasting methods in terms of computational and performance scalability. Our extensive ablation study reveals that all three spatial and temporal context data: elevation data, temporal covariates, and sensor observations (even if sparse), help Diff-Sparse make better forecasts.
Acknowledgments
This work was partially supported by the following grants: (i) NSF Grants CCF-1918656, OAC-1916805, RISE-2053013; (ii) AI Institute: Agricultural AI for Transforming Workforce and Decision Support (AgAID) award No. 2021-67021-35344; (iii) U.S. Department of Energy, through the Office of Advanced Scientific Computing Research’s “Data-Driven Decision Control for Complex Systems (DnC2S)” project. Pacific Northwest National Laboratory is operated by Battelle Memorial Institute for the U.S. Department of Energy under Contract No. DE-AC05-76RL01830.
Appendix A Related Work
@@
Hydrology Models: Traditionally, inundation has been simulated through hydro-dynamical physics-based 1D/2D dual drainage models. By simulating water flow in water bodies and coupling them with 1D or 2D hydrodynamic models, these models can predict inundation in surfaces. For example, the 2D model SOBEK (deltares2014sobek) consists of a 1D network with information about the initial water volume of rivers and 2D rectangular computational cells. By coupling these two parts, the model uses the Saint-Venant flow equation and 2D shallow water equation to calculate discharge between cells and determine the water level at each cell of the Digital Elevation Model (DEM). TELEMAC-2D (vu2015two) also solves 2D shallow water equation to simulate water flow. However, its primary purpose is to understand flow through water bodies. Therefore, this makes this model more suitable for coastal regions, contrary to SOBEK which is more suitable for urban areas. Polymorphic models like SCHISM (zhang2016seamless) can capture water flow at very high resolution, down to even a few meters due to their hybrid grinding system and localized coordinate system. Despite these models’ ability to model inundation accurately at fine resolution, the computation time associated with them makes them infeasible for real-time forecasting  (guo2021urban).
@@
Regression-based Models Early approaches in finding surrogates for hydrology models involved applying various regression-based methods. Chang et al. (chang2010clustering) first applied K-means clustering to pre-process data to find inundation control points. The authors then applied linear regression or neural networks to predict inundation for grids depending on whether they are linearly correlated with the control points or not. Bermúdez et al. (bermudez2019rapid) explored non-parametric regression approach based on least square support vector machine (SVM) as surrogate for 2D hydrodynamic models. Lee et al. (lee2021scenario) created an offline rainfall-runoff-inundation database by simulating hydrology models over multiple scenarios. Then, they trained a logistic regression-based approach based on the database to estimate the parameters to predict the risk of flooding for each region. Their goal was then to use the trained regressor to generate near real-time prediction of inundation probability given the runoff to the grid from simulated and realistic rainfall scenario.
@@
Deep Learning: DL methods are a popular choice in surrogating hydrodynamic models. Pan et al. (pan2011hybrid) applied a combined approach consisting of principal component analysis (PCA) and a feed-forward neural network. Their approach involved inundation prediction at various locations in the next timestep by looking at a fixed window of rainfall history of $N$ different rainfall gauges. First, they applied PCA on these various rainfall gauge histories to reduce the dimensionality before using them as input to the feed-forward neural network. Since such neural networks are expected to fail where data availability is limited, Xie et al. (xie2021artificial) explored the potential of a hybrid feed-forward neural network-based approach where they used two block-based neural networks along with the traditional point-based neural network and found superiority of the hybrid approach over other in the data-sparse region. Consequently, convolutional neural network (CNN) was explored (kabir2020deep) as a potential surrogate. After their model was trained to predict the output simulated from a 2D hydrology model, they used it to simulate two real-event flood scenarios in a region of UK. Since these models are not able to capture long-term temporal dependencies, researchers have recently explored (roy2023application) seq2seq LSTM model as surrogates to hydrology models. Such model was also applied in forecasting for multi-step future horizons. Aside from this, regression-based methods (zahura2022predicting) and hybrid ML-based methods (zanchetta2022hybrid) have been attempted as possible surrogates. However, these models cannot inherently take spatial context into account. The DCRNN method proposed by (li2017diffusion) for traffic forecasting is the state-of-the-art DL method for spatiotemporal point forecasting, where bidirectional random walk is used to capture spatial dependency and Seq2Seq architecture is used to capture temporal dependency.
@@
Diffusion Model The field of generative modeling has been dominated by the diffusion model for quite some time. Although it had primarily been studied for image synthesis (ho2020denoising; dhariwal2021diffusion; austin2021structured), its fascinating ability to use flexible conditioning to guide the generation process has been utilized in other domains like video generation (yang2023diffusion), prompt-based image generation (ramesh2022hierarchical), text-to-speech generation (yang2023diffsound). TimeGrad (rasul2021autoregressive) was the pioneering work that proposed the use of diffusion models for time-series forecasting tasks. They used temporal co-variates and embedded historical data as conditions for the diffusion model to denoise the prediction for the next time step. CSDI (tashiro2021csdi) does the same but instead of encoding with a Recurrent Neural Network (RNN) as done in TimeGrad, they leverage a transformer-based architecture. SSSDS4 (alcaraz2022diffusion) uses an S4 model (gu2021efficiently) to capture the encoding which is particularly useful in capturing long-term dependencies. Also, instead of performing diffusion along both temporal and input channel dimensions, they perform diffusion only along the temporal dimension. Other models like DPSD-CSPD (bilovs2023modeling), TDSTF (Chang2024) were later introduced. These models try to predict the noise added during the forward process of the diffusion model ($\epsilon$-parameterization). An alternative parameterization is to predict the ground-truth data ($x$-parameterization). Different from models mentioned above, $D^{3}$VAE (li2022generative) uses coupled diffusion process and bidirectional variational autoencoder to better forecast in longer horizon and improve interpretability. The State-of-the-Art Diffusion-based model DiffSTG (wen2023diffstg) devised a modified architecture for the reverse process of diffusion model combining UNet and GNN to capture temporal and spatial dependencies, respectively. However, both these methods scale poorly with the growth in number of variables. BayesNF (saad2024scalable) is the state-of-the-art DL spatiotemporal forecasting model, which is built using a Bayesian Neural Network. By assigning a prior distribution to the model parameters and adjusting the posterior based on observation, BayesNF essentially learns to map multivariate spatiotemporal points to a continuous real-valued field.
@@ Appendix B Denoising Diffusion Model
Let $\mathbf{x}^{0}\sim q_{\chi}(\mathbf{x}^{0})$ denote the multivariate training vector from some input space $\chi=\mathbb{R}^{D}$ and true distribution $q_{\chi}(\mathbf{x}^{0})$. In diffusion probabilistic models, we aim to approximate $q_{\chi}(\mathbf{x}^{0})$ by a probability density function $p_{\theta}(\mathbf{x}^{0})$ where $\theta$ denotes the set of trainable parameters. Diffusion probabilistic models (sohl2015deep; luo2022understanding) are a special class of Hierarchical Variational Autoencoders (kingma2016improved; sonderby2016ladder) with the form $p_{\theta}(\mathbf{x}^{0})=\int p_{\theta}(\mathbf{x}^{0:N})d\mathbf{x}^{1:N}$; here $\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{N}$ are latent variables. Three key properties of diffusion models are: $(i)$ all of the latent variables are assumed to have the dimension $D$ as $\mathbf{x}^{0}$, $(ii)$ the approximate posterior $q(\mathbf{x}^{1:N}|\mathbf{x}^{0})$,
@@
Here, we get (18) by applying Jensen’s inequality. We indirectly minimize the negative log-likelihood of the data, i.e., $\log p_{\theta}(\mathbf{x}^{0})$, by minimizing the right hand side of (19) . We can show that it is equal to:
@@
(ho2020denoising) showed that the forward process has the following property: we can sample $\mathbf{x}^{n}$ using $\mathbf{x}^{0}$ at any noise step $n$ in closed form . Let, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$. Then, we have:
@@
[…]
@@
[…]
@@
[…]
@@
$\mathbf{x}^{n-1}=\mu_{\theta}(\mathbf{x}^{n},n)+\sqrt{\Sigma_{\theta}}\ \mathbf{u}$ (31)
@@
Here, $\mathbf{u}\sim\mathcal{N}({\mathbf{0}},\mathbf{I})$ for $n\in[2,N]$ and $\mathbf{u}={\mathbf{0}}$ when $n=1$. $\mu_{\theta}(\mathbf{x}^{n},n)$ and $\Sigma_{\theta}$ are computed using (28) and (27), respectively.
Appendix C Evaluation Metric
We use two metrics to evaluate forecast quality. The first metric is Normalized Root Mean Squared Error (NRMSE). Let $\mathbf{x}$ be the set of all observations over all cells for all timesteps. Let $x_{ij,t}\in\mathbf{x}$ denote the observed inundation value of cell $(i,j)$ at time $t$ and let, $\tilde{x}_{ij,t}$ denote the corresponding average forecast value. Then:
In addition to NRMSE, we use an extended version of the Continuous Ranked Probability Score (CRPS) (winkler1996scoring) to quantify forecast uncertainty. The CRPS metric is calculated for each cell and each timestep individually. Let’s assume that we have $M$ forecasts $\hat{x}_{ij,t}^{1},\hat{x}_{ij,t}^{2},...,\hat{x}_{ij,t}^{M}$} for observation $x_{ij,t}$. First, the empirical cumulative distribution function (CDF) at a point $z$, $\hat{F}_{ij,t}^{M}(z)$ from these forecast values is defined as:
where $\mathbb{I}(y)$ is the indicator function that yields $1$ if condition $y$ holds, 0 otherwise. From this, CRPS between the empirical CDF $\hat{F}_{ij,t}^{M}$ and the observation $x_{ij,t}$ is calculated as
$\text{CRPS}(\hat{F}_{ij,t}^{M},x_{ij,t})=\int_{z}{\Big(\hat{F}_{ij,t}^{M}(z)-\mathbb{I}({x}_{ij,t}\leq z)\Big)^{2}dz}$ (34)
Here, $|x_{ij,t}|$ denotes the absolute value of $x_{ij,t}$.
Appendix D Additional Experiment Results
Varying Prediction Length
As Diff-Sparse forecasts the inundation levels in future timesteps auto-regressively, we can use it to forecast inundation levels for arbitrary number of future timesteps. However, this can result in error accumulation. To test this, we took the setting $(64^{2},10)$ and varied the prediction length with the values $1,4,12,20,28,36,44,52,60$. Figure 4 shows the values of the two performance metrics (NACRPS and NRMSE) for our test data, for each of these prediction lengths, where context length is set to 12 hours and sparsity level to $95\%$. We observe that Diff-Sparse performs best, in terms of both performance metrics, when prediction length is $1$. This is expected as Diff-Sparse is trained to make predictions for one future timestep. As the prediction length is increased, we observe that both NACRPS and NRMSE increases.
Figure 4: Line-plots showing the performance of Diff-Sparse for different prediction lengths with test dataset configuration $(64^{2},10)$, context length of 12 hours, and sparsity level of $95\%$, in terms of two metrics. The dots represent the mean value of the metrics over ten experiment runs with different seeds; the vertical lines represent standard deviation. We observe that Diff-Sparse performs best when prediction length is $1$. This is expected as Diff-Sparse is trained to make predictions for one future timestep. When making predictions for multiple timesteps auto-regressively, we observe that both NACRPS and NRMSE increase with increasing number of prediction timesteps.
Masking Strategy
In Diff-Sparse, we have presented a novel masking strategy (Algorithm 1, lines 1-1). In this strategy, we retain the inundation values in cells where sensors are placed (according to the sensor mask). For cells that do not have sensors, their inundation values are replaced with standard gaussian noise. We refer to this masking strategy as ”Noise Masking”. To show the efficacy of this masking strategy, we consider a baseline masking strategy as follows: retain the inundation values in cells where sensors are placed; cells that do not have sensors, their inundation values are replaced with the value 0. We refer to this baseline masking strategy as ”Zero Masking”.
Figure 5: Bar-plots showing the performance of Diff-Sparse (in terms of NACRPS and NRMSE) for two different masking strategies during training (Zero mask and Noise mask), with patch configuration $(64^{2},10)$ and $(64^{2},20)$ and sparsity level $95\%$. Context and prediction lengths are set to 12 hours. The height of the bars represent the mean value of the metrics over ten experiment runs with different seeds; the vertical black lines represent standard deviation. We observe that Diff-Sparse noise masking yields better result in terms of both performance metrics for the two patch configurations.
Figure 5 shows a comparison of the two masking strategies in terms of the resulting predictive performance of Diff-Sparse. We observe that, for the patch configurations $(64^{2},10)$ and $(64^{2},20)$ at $95\%$ sparsity level, noise masking yields better predictive performance than zero masking. Our hypothesis is, by placing random gaussian noise, noise masking provides an additional signal to the model to ignore the inundation values where sensors are not available.
Forecasting Time
To understand the efficacy of Diff-Sparse in real-time forecasting, we recorded its forecasting time. With context length of 12 hours, prediction length of 36 hours, and number of scenarios to be predicted set to $8$; Diff-Sparse takes $\sim 7$ seconds to forecast for a patch of size $64\times 64$, on a single GPU. Since each cell of the patch is $30m\times 30m$, area of a $64\times 64$ patch is $1.423$ sq. miles. The entire eastern shore of Virginia has a total area of $2105$ sq miles, which would require $\sim 1480$ patches of size $64\times 64$. On a single GPU, making 36 hour forecasts for all these patches would take $2.88$ hours using Diff-Sparse. However, we can parallelize the forecasting of different patches. For instance, if we have $8$ GPUs available, making $36$ hour forecast for the entire Eastern Shore of Virginia would take $2.87/8\approx 0.36$ hours or $\sim 22$ minutes.
Appendix E Inundation Scenario Queries